HUBBLE SPACE TELESCOPE DAILY REPORT #2062 PERIOD COVERED: 0000Z (UTC) 02/17/98 - 0000Z (UTC) 02/18/98 Daily Status Report as of 049/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 7621 (WF/PC-2 Cycle 7 Supplemental Darks PT1/3) The WF/PC-2 was used to make supplemental dark calibration images. This set of calibrations is intended to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed STIS/CCD 7926 (CCD Dark and Bias Monitor -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make several dark and bias calibration observations. This is to measure CCD dark current and bias rates for the two supported amplifier settings {gain = 1 and gain = 4} in order to gather data for making superdarks and superbiases. These observations will also track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed STIS/CCD 7711 (STIS CCD G750L Fringing Flats.) The Space Telescope Imaging Spectrograph (CCD) was used to investigate and monitor the STIS CCD flat-field fringing in the G750L grating at its nominal wavelength setting {7751A}. WAVECALs were also taken at the same gain and wavelength setting. Typical S/N = 100 for each integration, and gain = 4 on each. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed Five Sets of STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. The observations were completed as planned, and no problems were reported. 1.5 Completed WF/PC-2 7621 (WF/PC-2 Cycle 7 Supplemental Darks PT1/3) The WF/PC-2 was used to make supplemental dark calibration images. This set of calibrations is intended to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed STIS/CCD 7803 (Daily Darks to Update Acquisition Bad Pixel Table Part III) The Space Telescope Imaging Spectrograph (CCD) was used to make daily dark calibration updates. These daily darks are intended to identify hot pixels for update to the acquisition on-board FSW hot pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were completed as planned, and no anomalies were reported. 1.7 Completed NIC/2 7175 (H-2 Imaging of Photon-Dominated Regions {PDRs}) The Near Infrared Camera (Camera Two) was used to make imaging of the photon dominated regions in IC63. The Orion Bar feature southeast of the Trapezium is a well studied PDR with some very interesting chemical and physical features. Molecular hydrogen {H-2} emits strong near-infrared line emission. The observations were completed as planned, and no anomalies were reported. 1.8 Completed WF/PC-2 7625 (WF/PC-2 Cycle 7 Earth Flats) The WF/PC-2 was used to make a series of Earth flat calibration images. These flat fields will allow mapping of the OTA illumination pattern and will be used in conjunction with previous internal and external flats to generate new pipeline superflats. The observations were completed as planned, and no anomalies were reported. 1.9 Completed Two Sets of NIC/1/2/3 7899 (An IR Search For faint Companions to Pleiades Very Low-Mass Stars and Brown Dwarfs) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of ROQUE-11 and ROQUE-12. This is part of a infrared campaign to search for faint companions to Pleiades very low-mass stars and brown dwarfs. The Pleiades has been the best site for brown dwarf {BD} searches, with half of the presently "confirmed" BDs being Pleiades members. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed Three Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.11 Completed Two Sets of STIS/MA1 7651 (STIS MAMA Dispersion Solution Check) The Space Telescope Imaging Spectrograph (MAMA-1) was used to make MAMA dispersion solution checks. Wavelength dispersion solutions will be determined on a yearly basis as part of a long-term monitoring program. The observations were completed as scheduled, and no anomalies were reported. 1.12 Completed WF/PC-2/STIS/MA2 7410 (Deep UV Imaging of the Hubble Deep Field) The WF/PC-2 was used to make ultraviolet observations of Hubble Deep Field HDF-STIS-P. The observations will extend the searches for Lyman break galaxies to B~ 29, two magnitudes fainter than possible with the existing HDF data and more than four magnitudes fainter than the current limit of ground-based spectroscopy. The observations were executed as scheduled, and no anomalies were reported. 1.13 Completed NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel spectrographic observations of sky regions when another instrument is prime. The observations were completed as planned, and no anomalies were reported. 1.14 Completed NIC/1/2/3 7447 (The Cosmological Evolution of Quasar Host Galaxies) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of standard star PSF-STAR-2. Point spread function observations were taken to test the calibration of the cameras. The observations were executed as scheduled, and no anomalies were noted. 1.15 Completed NIC/3 7459 (The Age and Content of a {Proto}Galaxy Cluster at z=2.39) The Near Infrared Camera (Camera Three) was used to make observations of the proto-galaxy cluster Gal-Clust-1239. Our WF/PC-2 observations of the field around the radio galaxy 53W002 {z=2.39} have discovered a remarkable population of Lyman Alpha emission-line objects at the same redshift, with 18 identified in a single WF/PC-2 mosaic. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 6 The acquisition at 049/0230Z failed to fine lock on FGS-3 only. HSTAR 6347 was written. FOC proposal 6426 and WF/PC-2 proposal 6557, both occurring during the period of the next report, may have been affected. Scheduled Reacquisitions : 3 Successful : 2 Per HSTAR 6346, the re-acquisition scheduled for 049/070226Z failed, possibly affecting WF/PC-2 proposal 6557. This proposal will be described in tomorrow's report. 2.2 FHST Updates: Scheduled : 20 Successful : 20 2.3 Operations Notes: At 048/205556Z, NICMOS recorded an EDAC single bit error. There were no other error flags or additional status buffer messages. HSTAR 6345 was written. ROP NS-11 (to dump the EDAC RAM error log) was performed at 048/2212Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH